# Learning to Play Piano in the Real World

### v1

@@ I Introduction
Playing the piano requires humans to master contact-rich hand movements dictated by the timing and tone they intend to produce. This mastery is not learned quickly but through extensive practice, which requires humans to control their actions based on the haptic and auditory feedback received with each key pressed on the piano. In addition, human hands are an extraordinary research subject due to their unmatched dexterity, precision, and adaptability. These capabilities have evolved to perform various tasks, from delicate manipulations like sewing to powerful tasks like shot-put. This makes human-like hands a natural inspiration for robotic systems. In robotics, much attention has been paid to tasks such as in-hand object rotation, where the goal is to develop and optimize object manipulation through intricate movements [1, 15, 7, 8]. These studies often achieve efficient manipulation by exploiting unnatural and sometimes non-anthropomorphic motions to improve control and precision. Although such techniques yield impressive results, they often diverge from the natural movements of the human hand, which could limit their applicability to tasks that require more human-like actions. Playing a musical instrument like the piano is an exceptional example of human-like hand function.
@@
Piano playing demands a high level of motor control, coordination, and precision, all while closely aligning with the natural movements of human hands. This makes it an ideal scenario for exploring Sim2Real transfer, where the objective is to train an agent in simulation capable of performing in the real world. By focusing on playing piano with a human-like hand, this research aims to bridge the gap between the abstract, often unnatural solutions of robotic manipulation and the fluid, natural motions required to mirror the human approaches to complex tasks more closely.
In this work, we train a reinforcement learning policy in simulation that is applicable in the real world. The agent’s task is to play simple songs, such as the children’s song ”Twinkle Twinkle Little Star”, with a multi-finger Allegro hand on a real world piano. We define the observation space of the model that comprises the notes that need to be played, the current joint positions of the hand, and the subsequent keys pressed at that time step. Our work is based on the open-source training environment Robopianist-Suite [18]. However, since our primary objective is to deploy the learned policy in the real world, we made numerous changes at the system level. Moreover, due to the changes in the environment, the reward function originally used in the robopianist environment was no longer applicable. This pushed us to design a new reward function that is applicable to hands with an arbitrary number of fingers and does not depend on pre-known fingering data. Further, to bridge the Sim2Real gap, we utilized domain randomization of simulation parameters, such as joint dampening, friction coefficients, and spring stiffness. To apply the learned policy in the real world, we developed a system using ROS to control the robotics based on the actions of the agent.
@@ II Related work
Understanding the role of robotic systems in playing musical instruments goes a long way back to when highly sophisticated robots such as Takanishi’s Anthropomorphic Flutist Robot [12] that could mimic human respiratory systems were utilized to play the flute. Similarly, early works concerning robots playing the piano indicate the extensive effort made by researchers who built humanoid-like robots that leveraged precise finger-arm movement for rendering tunes [3]. Further works also explored the development of intelligent algorithms [6] and control systems [5] for anthropomorphic robots designed to play the piano. Although path paving, most of the works described heavily relied upon pre-programmed movements, limiting the robot’s ability to adjust to musical styles or improvise on existing scores. Therefore, we move away from traditional methods to a more learnable and robust methodology. Instead, we train the policy in the simulation to play a musical piece without any hard-coded or pre-programmed knowledge.
Moreover, playing the piano is a challenging benchmark for dexterity and the contributions in this field are very limited. The contribution of [11] focused on using inverse kinematics to play piano with Shadow hands in the real world, which does not utilize machine learning. In contrast, the contribution by Qian et al. [9], used imitation learning based on YouTube videos to play piano in a simulated environment. Research with robots and piano is novel, and initial works in simulation have demonstrated that learned agents can play simple songs and chords in a simulated environment [17, 18]. With the introduction of Robopianist [18], it was shown that by leveraging the reinforcement learning algorithms, an agent can be trained to play advanced music with two simulated Shadow hands. Furthermore, systematic reinforcement learning approaches in simulation elaborate on utilizing touch feedback from vision-based tactile sensors [4, 16], to teach a robot hand to play the piano [17]. Unlike most prior works, which focus on developing a paradigm to teach robot hands to play in simulation, we prioritize both teaching a robot hand to play the piano in simulation and translating this learned knowledge into a real world environment.
@@ III Learning to Play Piano in the Real World
Hardware Configuration. The hardware setup is shown in figure 2(b). We use a UFACTORY xArm7 to move the end effector along a specified axis parallel to the piano. This mimics the movement of the human forearm.
Attached to the end effector is a left-sided Allegro hand v4.0. It consists of 4 fingers with 4 joints per finger, leading to 16 joints in total. The thumb joints are not used in this contribution because they can barely reach the piano, leading to 12 active joints. Further, each of the fingertips is replaced with a 3D-printed finger of approximately the dimensions of a human finger. This is necessary because the default fingertips of the Allegro hand are wider than a key, which makes it impossible to play only one note at a time.
As a piano, we use the M-Audio Keystation 49e MIDI Keyboard, which contains four octaves. Every keypress and every release triggers a MIDI event, which is forwarded to the host computer. This allows the agent to observe the currently pressed keys, however, the exact degree of the keypresses is unknown.
@@
Simulation Environment. The environment is based on the robopianist environment [18], which was built using the Mujoco physics engine [13] with dm_control [14] as a wrapper for RL tasks. The scene of the simulation is displayed in Fig. 2(a). It consists of the Allegro hand with 3D printed fingertips and a model of the M-Audio Keystation piano. The hand is modeled analogues to the real world with 12 active joints, 4 per finger. To move the hand parallel to the piano, we introduced another sliding joint for horizontal movement. The piano itself consists of 49 keys and 4 octaves. We still use the synthesizer of the robopianist environment, which is based on fluidsynth, but the sustain pedal was removed since it is not part of the real world environment.
Markov-Decision-Process. The task is modeled as a stochastic, partial observable MDP defined by the tuple $(S,A,p,r,\gamma)$.
  • The state space $S$ and the transition probabilities $p$ are defined by the Mujoco physics engine. At each simulation step is the current state stored in the environment.physics.data object.
  • The action space $A$ consists of 13 continuous actions. One action per joint of the hand and one additional action for the wrist movement, parallel to the piano.
  • The observation space $O$ is 356 dimensional and is explained in more detail in the Table I. Note that the states of the keys are discrete. This makes playing the piano more demanding because the agent cannot hover with the fingers on the keys to observe the environment, as described in the robopianist [18]. Unfortunately, it is necessary because the MIDI standard used by the piano does not include partially pressed keys. Another contrast to the robopianist environment [18] is that the observation space does not include any fingering information. This is due to the Allegro hand having only 4 fingers, which makes regular ”human” fingering not applicable. The missing fingering information also has a negative impact on the exploration. For those reasons, a carefully chosen reward function is even more crucial to succeed in the exploration.
@@
Used joints of the Allegro Hand rad $12$
Slider for parallel movement m $1$
Currently pressed keys one-hot-encoded $49$
Currently targeted keys one-hot-encoded $49$
Targeted keys one-hot-encoded $5*49=245$
@@
Reward Design. The reward function consists of four parts
each rewards certain behavior of the agent. We will briefly discuss the purpose of each part in this section. However, a more detailed discussion is included in Appendix VI-A.
The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations that don’t lead to higher rewards.
The reward $r_{\text{hand\_position}}$ is designed to guide the hand toward targeted keys. This is achieved by reducing the average distance between the palm of the hand and the targeted keys.
  1. 1.
    Pressing no keys should be worse than pressing the wrong keys.
  2. 2.
Those requirements lead the exploration of the model towards pressing the correct keys without being ”afraid” of pressing the wrong keys.
The $r_{\text{sliding}}$ reward is designed to penalize fast sideways movement of the hand while keys are pressed. This avoids the local optimum where one finger is pressed statically while the wrist slider ”selects” the key. Further, the penalty is doubled if two adjacent keys are pressed, which is unavoidable while sliding sideways.
Evaluation Metrics. To quantify the success of the agent we apply the same metric as in the robopianist [18] contribution. At each timestep, the current states of all piano keys are compared to the target states, addressing two questions:
  • How reliable is the model at not pressing the keys it should not press? (Precision)
  • How reliable is the model at pressing the keys it should press? (Recall)
Both, the precision and recall score range from 00 to $1$. A score of $1$ indicates that the agent performed perfectly in all timesteps according in this metric. Both metrics are combined using the harmonic mean to calculate the F1 score
@@
We use the F1 score as our primary score to quantify the success of the agent. We also support the empirical statement made in the robopianist [18] that the F1 score represents the qualitative performance of the agent. Further, the F1 score is a common heuristic accuracy score in the audio information retrieval literature [10].
Figure 3: The diagram compares the three execution modes we introduced: A) In joint mirroring, the whole observation space is obtained from the simulated environment. B) In hybrid execution, only the pressed keys are based on the real world, while everything else is simulated. C) In real world execution, all observations are based on the real world.
Our Workflow. Our workflow is separated into two phases:
First, we start with training a policy exclusively in simulation. This policy can then be used to play piano in the real world with an execution mode. In the following sections, we will discuss both phases in more detail.
[…]
  • The height of the piano
  • The starting position of the hand
  • The dampening of the joints
  • The stiffness of the joints
  • The threshold of a key to be considered pressed
  • The stiffness of the springs holding the keys
  • The friction between the keys and the fingertips
[…]
Execution Modes. To evaluate the trained policy in the real world, we introduce three modes of execution, which are also displayed in Fig. 3:
  1. 1.
    Joint Mirroring: In this mode are all observations based on the simulation. The real world robot mirrors the joint positions of the simulation into the real world. This leads to a deterministic sequence of actions, but it is impossible to adapt to the environment.
  2. 2.
@@
  3. 3.
    Hybrid Execution: The observation space of the model is split into observations of the real world and observations of the simulation. This is possible because the simultaneous inference of the model and the simulation of the physics environment is possible in real time. For our purpose, the pressed keys are based on the real world, while the joint positions are based on the simulation.
Song Collection. In total, we validate our models on four different songs, each having a different characteristic:
  1. 1.
    Twinkle Twinkle Little Star: This song is a simple children’s song. Its most relevant characteristic is the big interval at the beginning of the song and the natural flow of the song. It further serves as an everyday live example.
  2. 2.
    C-Major-Scale: Having a scale in the reportorial is useful since it evaluates how the model changes from one key to another. For humans, it is intuitive to use all the fingers for playing the scale. We will discuss in the experiments whether our agent exploits the same strategy.
  3. 3.
    D-Major-Scale: The major difference between the C-Major scale and the D-Major scale is that the D-Major scale contains black keys as well. This allows us to observe the fingering technique of the agent while pressing black keys.
  4. 4.
    Chord Progression: In the first three songs, the agent only needs to press one key at a time. In this song we train on a series of four chords instead, which lets us evaluate the multi finger performance.
@@ IV Experimental Results
We provide videos showcasing real world runs for each of the experiments on our website: www.lasr.org/research/learning-to-play-piano.
In this section, we experimentally answer the following questions:
IV-A How does our model perform? (Fig. 4)
@@
Setup. To evaluate the performance of our approach, we trained four different songs. Each of those songs was trained twice on independent seeds, which led to eight models in total. Since every run in the real world differs, we executed each model three times. The figure shows the average and standard deviation of the precision, recall and the resulting F1 score for each song after using hybrid execution as introduced in Section 3. Other execution modes are evaluated in Section IV-B. As a reference for the Sim2Real gap is also the performance in simulation provided.
Figure 4: This diagram compares the performances of multiple songs in the real world with hybrid execution. For each song, we compare the precision, recall, and F1 score. As a reference is the F1 score, which was reached in the simulation, provided. The displayed data shows that although we observe a significant Sim2Real gap, we are able to play all four songs successfully. This indicates that our approach is generalizable and can be applied to other applications as well.
Results. All four songs have in common that the precision is higher than the recall score. A possible explanation is that optimizing the recall score is easier than optimizing the precision. When the robot hand presses a wrong key, then the agent will try to correct it. For the precision short, it is already sufficient to lift the finger since no wrong keys are pressed anymore. However, to improve the recall score, the agent needs to lift the finger, move it sideways, and press the correct key. This sequence of actions is significantly more demanding and takes more time. By the time this takes, a new key might be already targeted, resulting in skipping some keys.
Another explanation for the higher recall compared to the precision score is that maximizing the recall score is easier than maximizing the precision score. To maximize the precision, the agent needs to press at least all correct keys. But to maximize the recall, the agent can easily press no keys at all. This makes the precision, compared to pressing no keys at all, more demanding.
Another observation is that the Chord Progression performed the worst in the simulation, with an F1 score of $\approx 0.2$ less than all other songs. This might be because playing chords requires the robot to press multiple keys at once. Therefore, the robot needs to organize multiple fingers at the same time, which might require a bigger model than just using one finger with occasional finger switches.
However, the observed F1 score in the real world is significantly above the average, with almost no variation. This results in the smallest Sim2Real gap of this experiment. A possible reason is that each chord needs to be pressed for a longer duration than a single key in the other songs. This enables the robot to correct itself more easily, decreasing the Sim2Real gap.
Conclusion. Our models are able to play all four songs successfully. Although they have a significant Sim2Real gap, the experiment shows that our approach is generalizable to multiple songs. This demonstrates the high potential of hybrid execution for other contributions.
IV-B Which execution mode has the most potential? (Fig. 5)
@@
Setup. This experiment is an ablation study over our execution modes: joint mirroring, hybrid execution, and real world execution, which were introduced in Section III. No domain randomization was used for the joint mirroring and hybrid execution. For the real world execution we decided to use a moderate intensity of domain randomization with $C_{dr}=0.5$. The effect of domain randomization is further discussed in Section IV-C. Analog to the prior experiment, we trained each model configuration twice and executed each model 3 times per execution mode. Therefore, the figure shows the average and standard deviation of 6 runs for each mode.
Results.
Figure 5: The diagram compares our execution modes on multiple songs. The best performing mode is hybrid execution. It allows the agent to correct itself while the joints behave similarly to what the agent was trained on in simulation.
The diagram shows that the agent’s performance differs across the execution modes depending on the song. While joint mirroring performed among the best for Twinkle Twinkle Little Star and the Chord Progression, it performed worse than other modes for both scales. The real world execution, on the other side, performed well for the C-Major Scale but underperformed for all other songs. The hybrid execution was the best-performing mode for three out of four songs. The only exception is Twinkle Twinkle Little Star, where the hybrid execution mode was slightly behind joint mirroring. To understand the superiority of hybrid execution, we need to discuss the advantages and disadvantages of each execution mode separately.
When using joint mirroring, we don’t observe the real world, which makes it impossible to correct wrong keypresses. The advantage of this mode is that it does not need domain randomization, which reduces the resource requirements and the number of hyperparameters. The lack of adaptability is especially a problem for both scales since the model learned to slide sideways across the keys during the training. Unfortunately, this strategy is hard to execute in the real world since the finger slips differently every time. Since the agent does not observe the real world, it is also not able to correct those wrong movements.
The real world execution has the advantage of observing the whole state of the real world robot. This could allow the robot to react in the best way since no information is hidden. Unfortunately, the additional information also requires a more robust model, which can be reached by utilizing domain randomization as shown in Section IV-C. However, domain randomization makes the training environment significantly more demanding. In our experiments, the model was not able to outperform the other execution modes.
The hybrid execution tries to combine both advantages. Most of the observations are still generated in simulation, which makes it easier for the agent to apply the learned strategies. However, by providing observations of the actually pressed keys, the robot is able to correct its behavior. This way, the agent behaves similarly to joint mirroring when the pressed keys match the currently pressed keys in the simulation. However, as soon as the simulation differs from the real world, the agent is able to correct itself. This is possible while not using any domain randomization at all, which makes the process even more favorable.
@@ Results.
Conclusion. We showed that hybrid execution is the most promising execution mode for playing piano. It allows the agent to correct itself while being able to use the joints in the same way as they behaved during training. This makes hybrid execution modes especially interesting for future work with the goal of achieving human-like proficiency.
IV-C How does the intensity of domain randomization affect the real world performance? (Fig. 6, Fig. 7)

### v2

@@ I Introduction
Playing the piano requires humans to master contact-rich hand movements dictated by the timing and tone they intend to produce. This mastery is not learned quickly but through extensive practice, which requires humans to control their actions based on the haptic and auditory feedback received with each key pressed on the piano. In addition, human hands are an extraordinary research subject due to their unmatched dexterity, precision, and adaptability. These capabilities have evolved to perform various tasks, from delicate manipulations like sewing to powerful tasks like shot-put. This makes human-like hands a natural inspiration for robotic systems. In robotics, much attention has been paid to tasks such as in-hand object rotation, where the goal is to develop and optimize object manipulation through intricate movements [1, 14, 7, 8]. These studies often achieve efficient manipulation by exploiting unnatural and sometimes non-anthropomorphic motions to improve control and precision. Although such techniques yield impressive results, they often diverge from the natural movements of the human hand, which could limit their applicability to tasks that require more human-like actions.
@@
Playing a musical instrument like the piano is an exceptional example of human-like hand function. Piano playing demands a high level of motor control, coordination, and precision, all while closely aligning with the natural movements of human hands. This makes it an ideal scenario for exploring a Simulation to Real World (Sim2Real) transfer, where the objective is to train an agent in simulation capable of performing in the real world. Our work focuses on the challenges of this transfer. We provide interesting insights into the nature of playing piano and actively encourage researchers to participate in this field of study. We believe that by mastering piano playing in the real world, we can bridge the gap between the abstract, often unnatural solutions of robotic manipulation and the fluid, natural motions required to mirror the human approaches to complex tasks more closely.
In this work, we train a reinforcement learning (RL) policy in simulation that is applicable in the real world. The agent’s task is to play simple songs, such as the children’s song ”Twinkle Twinkle Little Star”, with a multi-finger Allegro hand on a real world piano. Our work is based on the open-source training environment RoboPianist-Suite [17]. However, since our primary objective is to deploy the learned policy in the real world, we made numerous changes at the system level. Moreover, due to the changes in the environment, the reward function originally used in the RoboPianist environment was no longer applicable. This pushed us to design a new reward function that is applicable to hands with an arbitrary number of fingers and does not depend on pre-known fingering annotations. To bridge the Sim2Real gap, we utilized domain randomization (DR) of simulation parameters, such as joint dampening, friction coefficients, and spring stiffness.
This work contributes in multiple ways:
  • By providing a proof-of-concept, we want to encourage the community to adopt piano playing as a compelling benchmark towards human-level manipulation.
  • By analyzing three separate execution modes, we gain a deep understanding of the challenges for playing piano in the real world. These observations can be used to develop more capable models.
  • Our model showcases the success with common RL techniques. The results can be used as a baseline to evaluate alternative approaches.
  • Our codebase is open-sourced, which provides researchers a headstart for implementing and verifying new approaches for playing piano in the real world.
@@ II Related work
Understanding the role of robotic systems in playing musical instruments goes a long way back to when highly sophisticated robots such as Takanishi’s Anthropomorphic Flutist Robot [11] that could mimic human respiratory systems were utilized to play the flute. Similarly, early works concerning robots playing the piano indicate the extensive effort made by researchers who built humanoid-like robots that leveraged precise finger-arm movement for rendering tunes [3]. Further works also explored the development of intelligent algorithms [6] and control systems [5] for anthropomorphic robots designed to play the piano. Although path paving, most of the works described heavily relied upon pre-programmed movements, limiting the robot’s ability to adjust to musical styles or improvise on existing scores. Therefore, we move away from traditional methods to a more learnable and robust methodology. Instead, we train the policy in the simulation to play a musical piece without any hard-coded or pre-programmed knowledge, including fingering annotations.
Moreover, playing the piano is a challenging benchmark for dexterity and the contributions in this field are very limited. The contribution of [10] focused on using inverse kinematics to play piano with Shadow hands in the real world, which does not utilize machine learning. In contrast, the contribution by Qian et al. [9], used imitation learning based on YouTube videos to play piano in a simulated environment. Research with robots and piano is novel, and initial works in simulation have demonstrated that learned agents can play simple songs and chords in a simulated environment [16]. Those RL approaches in simulation elaborate on utilizing touch feedback from vision-based tactile sensors [4, 15], to teach a robot hand to play the piano [16]. With the introduction of RoboPianist [17], it was shown that by leveraging the RL algorithms, an agent can be trained to play advanced music with two simulated Shadow hands. Further, the RoboPianist environment was used to build a dataset containing $\approx$1 million trajectories of bimanual piano playing in simulation [18], which can be used to train an imitation learning model. Unlike most prior works, which focus on developing a paradigm to teach robot hands to play in simulation, we focus on translating this learned knowledge into a real world environment.
@@ III Learning to Play Piano in the Real World
Hardware Configuration. The hardware setup is shown in figure 2(b). We use a UFACTORY xArm7 to move the end effector along a specified axis parallel to the piano. This mimics the movement of the human forearm. Attached to the end effector is a left-sided Allegro hand v4.0. It consists of 4 fingers with 4 joints per finger, resulting in a total of 16 joints. The thumb joints are not used in this contribution because they can barely reach the piano, leading to 12 active joints. Each of the fingertips is replaced with a 3D-printed finger of approximately the dimensions of a human finger. This is necessary because the default fingertips of the Allegro hand are wider than a key, which makes it impossible to play only one note at a time. Our piano is an M-Audio Keystation 49e MIDI Keyboard, which contains four octaves. Every keypress and every release triggers a MIDI event, which is forwarded to the host computer.
@@
Simulation Environment. The environment is based on the RoboPianist environment [17], which was built using the Mujoco physics engine [12] with dm_control [13] as a wrapper for RL tasks. The scene of the simulation is displayed in Fig. 2(a). It consists of the Allegro hand with 3D-printed fingertips and a model of the M-Audio Keystation piano. The hand is modeled analogues to the real world with 12 active joints, 4 per finger. To move the hand parallel to the piano, we introduced another sliding joint for horizontal movement. The piano itself consists of 49 keys and 4 octaves.
@@
Used joints of the Allegro Hand rad 1212
Slider for parallel movement m 11
Pressed keys one-hot-encoded 4949
Intended keys one-hot-encoded 4949
Intended keys of future timesteps one-hot-encoded 549=2455*49=245
@@
Markov-Decision-Process. The action space consists of 13 continuous actions. One action per joint of the hand and one additional action for the wrist movement, parallel to the piano. The observation space is 356-dimensional and is shown in Table I. Note that the observations of the keys are discrete. This makes playing the piano more demanding because the agent cannot hover with the fingers on the keys to observe the environment, as described in the RoboPianist [17]. Unfortunately, it is necessary because the MIDI standard used by the piano does not include partially pressed keys. Another contrast to the RoboPianist environment [17] is that the observation space does not include any fingering annotations. This is due to the Allegro hand having only 4 fingers, which makes regular ”human” fingering not applicable. The missing fingering annotations also has a negative impact on the exploration. For those reasons, a carefully chosen reward function is even more crucial to succeed in the exploration.
Reward Design. We will briefly discuss each component of our reward function $r_{\text{total}}=r_{\text{energy}}+r_{\text{hand\_position}}+r_{\text{keypress}}+r_{\text{sliding}}$ in this section. However, a more detailed discussion is included in Appendix VI-A. The $r_{\text{energy}}$ reward penalizes fast movements that need a lot of force. Further, it avoids oscillations that don’t lead to higher rewards. The $r_{\text{sliding}}$ reward is designed to penalize fast sideways movement of the hand while keys are pressed to avoid damaging the robot. The reward $r_{\text{hand\_position}}$ is designed to guide the wrist towards intended keys, while $r_{\text{keypress}}$ encourages the agent to press the correct keys.
Evaluation Metrics. To quantify the success of the agent, we use the same metric as in the RoboPianist [17]. At each timestep, the current states of all piano keys are compared to the target states, addressing two questions:
  • ”How reliable is the model at not pressing the keys it should not press?” (Precision) and
  • ”How reliable is the model at pressing the keys it should press?” (Recall).
Both, the precision and recall score, range, from 0 to $1$. A score of $1$ indicates that the agent performed perfectly in this metric at all timesteps. Both metrics are combined using the harmonic mean to calculate the F1 score.
@@
We chose the F1 score as our primary score to quantify the success of the agent. Our driving argument is that we support the empirical statement, which was already made in the RoboPianist [17], that the F1 score represents the qualitative performance of the agent. Further, using the same metric across contributions makes the results more comparable and increases the transparency.
Figure 3: The diagram compares the three execution modes: A) In joint mirroring, the whole observation space is obtained from the simulated environment. B) In hybrid execution, only the pressed keys are based on the real world, while everything else is simulated. C) In real world execution, all observations are based on the real world.
Model Training. Our model consists of three hidden layers with 256 nodes each. The training uses DroQ [2], a soft actor-critic algorithm and runs for $1e6$ episodes. To reduce the Sim2Real gap, we utilized domain randomization (DR) of the height of the piano, the starting position of the hand, the dampening of the joints, the stiffness of the joints, the threshold of a key to be considered pressed, the stiffness of the springs holding the keys and the friction between the keys and the fingertips. Each simulation timestep is $0.005\text{\,}\mathrm{s}$. The current state is observed every 10th iteration, which leads to a control frequency of the policy of $20\text{\,}\mathrm{Hz}$.
Execution Modes. To deploy the trained policy in the real world, we introduce three modes of execution, which are also displayed in Fig. 3:
Joint Mirroring: In this mode, all observations are based on the simulation. The real world robot mirrors the joint positions of the simulation into the real world. This leads to a deterministic sequence of actions, but it is impossible to adapt to the environment.
@@
Hybrid Execution: The observation space of the model is split into observations of the real world and observations of the simulation. This enables the model to be deployed in the real world while being less robust compared to the real world execution. Hybrid Execution comes with the risk of a potentially diverging simulation, which needs to be tested carefully. In this work, the pressed keys are based on the real world, while the joint positions are based on the simulation.
Song Collection. We validate our models on multiple songs with different characteristics:
Twinkle Twinkle Little Star: This song is a simple children’s song. Its most relevant characteristic is the big interval at the beginning of the song and the memorable rhythm of the song. It further serves as an everyday life example.
C-Major-Scale: Having a scale in the repertoire is useful since it evaluates how the model changes from one key to another. For humans, it is intuitive to use all the fingers for playing the scale. We will discuss in the experiments whether our agent exploits the same strategy.
D-Major-Scale: The major difference between the C-Major scale and the D-Major scale is that the D-Major scale contains black keys as well. This allows us to observe the fingering technique of the agent while pressing the black keys.
Chord Progression: In the first three songs, the agent only needs to press one key at a time. In this song, we train on a series of four chords instead, which lets us observe the multi-finger performance.
@@ IV Experimental Results
IV-A How does our model perform? (Fig. 4(a))
@@
Setup. To evaluate the performance of our approach, we trained on four different songs. Each of those songs was trained twice on independent seeds, which led to eight models in total. Since every run in the real world differs, we executed each model three times. The figure shows the average and standard deviation of the precision, recall and the resulting F1 score for each song after using hybrid execution as introduced in Section 3. Other execution modes are evaluated in Section IV-B. As a reference for the Sim2Real gap is also the performance in simulation provided.
(a) Comparison of several songs in the real world using hybrid execution.
(b) Comparison of the three execution modes across multiple songs.
Figure 4: As a reference is in both diagrams the F1 score reached in the simulation is provided. The results show that, with our setup, hybrid execution is the best-performing execution mode. This indicates that the simulation does not significantly diverge from the real world, which enables us to improve upon the performance of the execution mode real world execution. The results also show the potential of hybrid execution in settings where it is hard to achieve a model robust enough for a complete real world execution.
Results. All four songs have in common that the precision is higher than the recall score. A possible explanation is that recovering from a wrong keypress consists of multiple steps. For the precision, just lifting the finger is sufficient since no wrong keys are pressed anymore. However, to improve the recall score, the agent needs to lift the finger, move it sideways, and press the correct key. This sequence of actions is significantly more demanding and takes more time. By the time this takes, a new key might be already targeted, resulting in skipping some keys. Another explanation for the higher recall compared to the precision score is that maximizing the recall score is a simple task compared to maximizing the precision score. To maximize the precision, the agent needs to press at least all the correct keys. But to maximize the recall, the agent can easily press no keys at all. Another observation is that the Chord Progression performed the worst in the simulation, with an F1 score of $\approx 0.2$ less than all other songs. This might be because playing chords requires the robot to press multiple keys at once. Therefore, the robot needs to organize multiple fingers at the same time, which might require a bigger model than just using one finger with occasional finger switches. Further, playing with multiple fingers simultaneously complicates the exploration of fingering techniques. However, the observed F1 score of the Chord Progression is significantly above the average, with almost no variation. This results in the smallest Sim2Real gap of this experiment. A possible reason is that each chord needs to be pressed for a longer duration than a single key in the other songs. This enables the robot to correct itself more easily, decreasing the Sim2Real gap.
Conclusion. In this experiment, we demonstrated that it is possible to train a piano-playing model using common RL techniques that can be deployed using hybrid execution. Our models are able to play all four songs successfully in simulation and in the real world. This demonstrates effectively that playing piano in the real world with a learned model is within the reach of our current capabilities.
IV-B Which execution mode has the most potential? (Fig. 4(b))
@@
Setup. In this experiment we perform an ablation study over the execution modes introduced in Section III: joint mirroring, hybrid execution, and real world execution. No DR was used for the joint mirroring and hybrid execution. For the real world execution we decided to use a moderate intensity of DR with $C_{dr}=0.5$. The effect of DR is further discussed in Section IV-C. Analog to the prior experiment, we trained each model configuration twice and executed each model 3 times per execution mode. The figure shows the average and standard deviation of 6 runs for each mode.
Results. Results show that the agent’s performance differs across the execution modes depending on the song. While joint mirroring performed among the best for Twinkle Twinkle Little Star and the Chord Progression, it performed worse than other modes for both scales. The real world execution, on the other side, performed well for the C-Major Scale but underperformed for all other songs. The hybrid execution was the best-performing mode for three out of four songs. The only exception is Twinkle Twinkle Little Star, where the hybrid execution mode was slightly behind joint mirroring. To understand the superiority of hybrid execution, we need to discuss the advantages and disadvantages of each execution mode separately. When using joint mirroring, we don’t observe the real world, which makes it impossible to correct wrong keypresses. Further, the simulation might diverge from the real world which would reduce the performance. The advantage of this mode is that it does not need DR, which reduces the resource requirements and the number of hyperparameters. The lack of adaptability is especially a problem for both scales since the model learned to slide sideways across the keys during the training. Unfortunately, this strategy is hard to execute in the real world since the finger slips differently every time. The real world execution has the advantage of observing the whole state of the real world robot. This ideally allows the robot to react in the best way possible since no information is hidden. Unfortunately, the additional information also requires a more robust model, which can be reached by utilizing DR as shown in Section IV-C. However, DR makes the training environment significantly more demanding. In our experiments, the real world execution mode was not able to outperform the other execution modes. The hybrid execution combines both modes. Most of the observations are still generated in simulation, which makes it easier for the agent to transfer the learned knowledge into the real world. However, by providing observations of the actually pressed keys, the robot is able to correct its behavior. This way, the agent behaves similarly to joint mirroring when the pressed keys match the currently pressed keys in the simulation. However, as soon as the simulation differs from the real world, the agent is able to correct itself. This is possible while not using any DR at all, which makes the process even more favorable.
@@
Conclusion. We showed that hybrid execution performed the best in our setup. It allows the agent to correct itself while being able to use the joints in the same way as they behaved during training. This makes hybrid execution modes especially interesting for future work with the goal of achieving human-like proficiency in complicated settings, where it is hard to deploy a model entirely into the real world.
IV-C How does the intensity of DR affect the real world performance? (Fig. 5(a), Fig. 5(b))